A roket starts from rest and accelerates with 2g.after height h engine is stopped.what is total height reached [ ans:3h]

Grade:11

2 Answers

Arun
3 years ago
Dear student
 
Please attach the image. As the language of question is ambiguous.
 
Hope you understand
 
Thanks and regards
Vikas TU
3 years ago
Dear student 
Accln= 19.6 m/s time t =5sec
Hmax = u square /2g
for initial velocity.
a=u/t i.e. 19.6=u/5
:. u=19.6×5=98m/s
:. Hmax=98 square/2(10)
9604/20=480.2m